OpenStudy (elifbestegul):

a metal requires energy of at least 4.0 eV to emit an electron from its surface determine the longest wavelength of light that will emit electrons

OpenStudy (irishboy123):

\(E = \dfrac{h c}{\lambda} \ge 4 eV\) \(\lambda \le \dfrac{h c}{4 eV} \) Use h, Plank's constant, in correct units. 

OpenStudy (osprey):

Photo electric effect (2). Shamrock's got the numbers it seems.

OpenStudy (irishboy123):

@osprey yes!! i saw your work earlier gonna have a stab at that gluon Spesh Rel thing later, mabe's. just need to get my blunderbuss out to dust off the cobwebs :-))

OpenStudy (osprey):

@IrishBoy123 My spec rel is about as bad as my ... I've been ferreting around, and my ferret's told me that, as usual, something's not being said somewhere as should be said. I'm beginning to wonder if you can actually get the initial muon speed out of the numbers posted. Another problem, of course, as with all these posts is time. how long to give any one q when it's just a one off q. Better if someone actually started some sort of "continuity thread" so that the time/mental effort investment was more cost effective. Pie in the sky ? Where's your blunderbuss ??????????

OpenStudy (irishboy123):

well if it's a muon, as opposed to a gluon ,...... LOL!!!! i'll need a flamethrower now, maybe @elifbestegul sorry for the banter on your thread, are you done now?!

OpenStudy (elifbestegul):

@IrishBoy123 no problem at all :) yeah I have done it, thank you so much